# Learning 3-Manifold Triangulations: Pachner Graph Network Analysis & Machine Learning IsoSigs

### v1

@@ 1 Introduction
Machine learning covers a broad range of techniques for computational statistics, and the subfield itself can be split into three core subsubfields, which each have already seen compelling successes in mathematics and mathematical physics. These are: supervised learning [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], unsupervised learning [26, 27, 28], and reinforcement learning [29, 30, 31, 32, 33, 34]. This work focuses on techniques from the first of these subsubfields, in particular using supervised learning methods to classify and study $3$-manifolds.
@@ 2.1 Triangulation IsoSigs
Each triangulation can be encoded by numbering the tetrahedra along with their vertices, indicating which pairs of faces of two tetrahedra of the triangulation are glued and how (there are $6$ possible identifications for each pair of faces). All these data can be encoded in a single string of characters defined in [48], called the isomorphism signature (IsoSig) of the triangulation. The following holds:
@@
It should be noted though, that a $3$-manifold admits infinitely many different (non-isomorphic) triangulations, which are all related to each other via the so-called bistellar or Pachner moves [45]. When applying a Pachner move, in general the IsoSig of a triangulation changes (for most moves even the number of tetrahedra changes). One can then build a graph $\mathcal{P}$ whose nodes are the isomorphism types of triangulations of a fixed manifold and whose edges represent that the two triangulations can be transformed into one another via a possible Pachner move [^4] Note that there may be (likely rare) scenarios where two different moves can transform an input triangulation into the same output triangulation, in these cases we still only connect the respective nodes by a single edge.. By the above cited fundamental result this graph is connected [45]. Before proceeding we remind the reader that these moves are of four types: $2-3$, $3-2$, $1-4$ and $4-1$ moves where for instance a $2-3$ move consists in exchanging two tetrahedra with a suitable combination of three tetrahedra and a $1-4$-move consists in subdividing a tetrahedron into four tetrahedra by coning from its barycenter to its vertices. The moves $3-2$ and $4-1$ are the inverses of respectively the $2-3$ and the $1-4$-moves. Note that the number of vertices of the triangulation does not change under $2-3$ and $3-2$ moves, but does change under $1-4$ and $4-1$ moves, exactly by one unit. Diagrammatics for these moves are shown in Figure 1.
@@
2.2 Algorithms for $3$-manifolds
@@ 3 Data Generation & Analysis
To examine and learn the 3-manifold triangulation isomorphism via IsoSigs, we select a variety of 3-manifolds to focus on, in particular those which span a range of interesting properties. Eight 3-manifolds were selected for these investigations, they include: the 3-sphere $S^{3}$; the product of the 2-sphere and the 1-sphere $S^{2}\times S^{1}$; 3-dimensional real projective space $\mathbb{RP}^{3}$, the lens space $L(7,1)$, the lens space $L(7,2)$; the 3-torus $T^{3}$; the Poincaré homology sphere $PHS$; and the smallest closed hyperbolic manifold $H_{SC}$. The 3-manifold abbreviations, along with the initial IsoSigs used to generate the datasets of equivalent triangulation IsoSigs are given in Table 1. Of these manifolds $S^{2}\times S^{1}$ is the only one not admitting a constant curvature isotropic complete Riemannian metric (we shall say for short a ‘non-isotropic manifold’), whilst the others are split amongst their curvatures, with $(S^{3},\mathbb{RP}^{3},L(7,1),L(7,2),PHS)$ positive curvature, $T^{3}$ zero curvature, and $H_{SC}$ negative curvature. In addition they have properties as listed in Table 2.
@@
For each 3-manifold’s initial triangulation and IsoSig, a dataset was generated of equivalent triangulations (each represented with their respective IsoSig) from the manifold’s Pachner graph computed exhaustively up to as high a depth (i.e. integer number of moves away from the initial IsoSig) as feasible within memory limits. The Pachner graphs were generated with moves $\{2-3,3-2\}$ (with further functionality in the code to consider $\{1-4,4-1\}$ moves). Initial investigations showed that ML architectures relied heavily on the IsoSig length for distinguishing manifolds, with performance measures $>0.99$ for near-perfect learning. Since the IsoSig length is heavily correlated with the number of tetrahedra, when we initiate generation using triangulations with a similar (e.g. small) number of tetrahedra then the graphs will display the same IsoSig lengths throughout, however the frequency distribution of the observed lengths varies, and this is what the architectures can leverage to provide efficient estimates of the considered manifold (note the final depth often all have the same length and dominate these distributions). For the manifolds considered the IsoSig lengths observed in the deep Pachner graph generation were: $(9,11,14,17,19,22,25,27,30,33,35,38,41)$. The respective distributions are shown in Figure 2.
@@ 3.1.1 Orientable Cusped Census
Figure 6: The distribution of the sizes of the Pachner graphs (i.e. number of nodes) as a function of (a) the volume and of (b) the length of the shortest geodesic. In blue the points corresponding to one cusped manifolds, in red those corresponding to $2$-cusped manifolds, and in green the remaining ones.
@@
We also analysed the size of the Pachner graphs as a function of the hyperbolic volume of the manifolds and their number of cusps (which varies from $1$ to $3$ in our dataset). The result is shown in Figure 6(a). Looking at the figure one realises that there are ‘peaks’ in the sizes of the Pachner graphs for one cusped manifolds and that these peaks correspond to the first red dots, namely the lowest volume $2$-cusped manifolds. This is no random phenomenon. The blue dot with highest size of the Pachner graph and volume $<3.63$ turns out to be “v0117” (in snappy’s nomenclature, IsoSig: hLALPkbcbefggghxwnxibk) and drilling out its shortest geodesic one gets “m129” (IsoSig: eLPkbdcddhgggb) which is the second lowest volume manifold in snappy’s census with number of cusps $2$ (it corresponds to one of the two leftmost red dots in the figure). This leads us to think that the size of the Pachner graph could be inversely related to the length of the shortest geodesic in the hyperbolic manifold which, by Thurston-Jorgensen’s theorem on hyperbolic Dehn fillings corresponds to Dehn fillings of manifolds with one more cusp with very large filling slopes. In order to check this we plotted in Figure 6(b), the sizes of the Pachner graphs with respect to the length $\ell$ of the shortest geodesic in the cusped manifold and the plot clearly shows that such number is bounded below by a function of $1/\ell$. Furthermore, a closer look reveals that the dots come in families, corresponding to which manifold is obtained by drilling the shortest geodesic. This is exemplified in Figures 7(a) and 7(b) where the same analysis as before is restricted to the manifolds whose drilling is a fixed one, here “m129” (chosen just as an example). All this seems to corroborate the following
@@
The growth rate of the Pachner graph of a cusped hyperbolic $3$-manifold $M$ is correlated to the inverse of the length of the shortest geodesic in $M$.
@@ 4.1 Differentiating Manifolds
The NN architecture is the prototypical machine learning model, it is very adaptable, and used ubiquitously with great success throughout many fields. Any NN is made up of sets of neurons organised into layers, where each neuron takes in a vector of inputs, and outputs a single number. The neuron acts on the input vector x with a linear function, then a non-linear ‘activation’ function [^10] Here we use the leaky-ReLU activation function, which has significant speed advantages due to its simplicity, and effectively makes the training process the fitting of a piecewise linear function., such that: $\textbf{x}\mapsto act(\textbf{w}\cdot\textbf{x}+b)$ for the trainable parameters of weights vector w and bias scalar $b$.
@@ 4.1.1 Neural Networks
We performed an in-depth analysis of the gradient saliency values used by our NN to distinguish $L(7,1)$ vs $L(7,2)$. A trained NN represents a function from inputs to outputs, in the case of these problems this is an input of a single IsoSig, one-hot encoded from length 30 in a length 64 alphabet to a length 1920 vector and an output of a length 2 vector with each entry evaluating in the range $[0,1]$ (as a probability distribution over the 2 classes) whose larger value indicates the predicted manifold the input IsoSig triangulation represents.
@@
Further examining the actual saliency values (represented by colours in Figure 12), averaged over all of the cross-validation runs as well as each runs test dataset, we can plot a histogram of the average saliency size, as shown in Figure 13(a). This image shows a partition of values with a selection having normalised gradients $>10^{-4}$ and the majority having values $<10^{-6}$. There is therefore a clear selection of ‘important’ inputs and ‘non-important’ ones. Taking the input indices corresponding to the larger ‘important’ saliency values $>10^{-4}$, these can be sorted into what character in the alphabet they correspond to (as plotted in a histogram in Figure 13(b)), or into what character in the length 30 IsoSig they correspond to (as plotted in a histogram in Figure 13(c)).
@@
Figure 13(b) shows that the 363 ‘important’ inputs correspond to only a subset of the full alphabet – predominantly the beginning of this alphabet with a few letters deeper in. Equivalently, these same important input features have much higher frequencies towards the end of the IsoSig as shown in Figure 13(c). From this saliency analysis the NNs are hence relying most on the final entries of the IsoSigs (the permutation sequence) to determine the manifold, and the value of these entries becomes discriminant. Additionally the occurrence of specific letters (like h or i with the highest frequency in Figure 13(b)) can significantly help the architectures confidence in classification.

### v2

@@ 1 Introduction
Machine learning covers a broad range of techniques for computational statistics, and the subfield itself can be split into three core subsubfields, which each have already seen compelling successes in mathematics and mathematical physics. These are: supervised learning [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] (introduced in [27]), unsupervised learning [28, 29] (introduced in [30]), and reinforcement learning [31, 32, 33, 34, 35, 36] (introduced in [37]); with notable introduction to string theory in [38, 39, 40, 41]. This work focuses on techniques from the first of these subsubfields, in particular using supervised learning methods to classify and study $3$-manifolds.
@@ 2.1 Triangulation IsoSigs
Each triangulation can be encoded by numbering the tetrahedra along with their vertices, indicating which pairs of faces of two tetrahedra of the triangulation are glued and how (there are $6$ possible identifications for each pair of faces). All these data can be encoded in a single string of characters defined in [55], called the isomorphism signature (IsoSig) of the triangulation.
IsoSigs, are an exceptionally compact representation method for 3-manifold triangulations, introduced in detail with examples in Appendix A, they are formed of a sequence of letters from the alphabet {a,b,...,z,A,B,...Z,0,1,...,9,+,-}, encoding numbers in base 64 which enumerate in turn: the number of tetrahedra in the triangulation, the destination sequence of face gluings, the type sequence listing boundaries, and the permutation sequence dictating the vertex matchings of the face gluings. The encoding relies on many tricks for removing redundancies to make the IsoSig encoding especially efficient compared to traditional methods of listing each of these components separately. Importantly for the IsoSig representation, the following holds:
@@
Where two triangulations are isomorphic if they are combinatorically equivalent, such that some relabelling of tetrahedra, faces, edges, and vertices makes them the same. It should be noted though, that a $3$-manifold admits infinitely many different (non-isomorphic) triangulations, which are all related to each other via the so-called bistellar or Pachner moves [52]. When applying a Pachner move, in general the IsoSig of a triangulation changes (for most moves even the number of tetrahedra changes). One can then build a graph $\mathcal{P}$ whose nodes are the isomorphism types of triangulations of a fixed manifold and whose edges represent that the two triangulations can be transformed into one another via a possible Pachner move [^4] Note that there may be (likely rare) scenarios where two different moves can transform an input triangulation into the same output triangulation, in these cases we still only connect the respective nodes by a single edge.. By the above cited fundamental result this graph is connected [52]. Before proceeding we remind the reader that these moves are of four types: $2-3$, $3-2$, $1-4$ and $4-1$ moves where for instance a $2-3$ move consists in exchanging two tetrahedra with a suitable combination of three tetrahedra and a $1-4$-move consists in subdividing a tetrahedron into four tetrahedra by coning from its barycenter to its vertices. The moves $3-2$ and $4-1$ are the inverses of respectively the $2-3$ and the $1-4$-moves. Note that the number of vertices of the triangulation does not change under $2-3$ and $3-2$ moves, but does change under $1-4$ and $4-1$ moves, exactly by one unit. Diagrammatics for these moves are shown in Figure 1.
@@
2.2 Existing Results Surrounding Pachner Graph Paths
@@
Since a theoretical algorithm exists, but is likely exceptionally slow and computationally expensive, there is strong motivation for discovery of other methods for identifying equivalence of triangulations (that they represent the same 3-manifold and are in the same Pachner graph). Existing results show promise for identifying paths to triangulations with fewer tetrahedra, but not between triangulations with general numbers of tetrahedra. It is hence well motivated to tackle this broader general equivalence problem, where this work initiates the application of ML methods for this.
@@ 3 Data Generation & Analysis
To examine and learn the 3-manifold triangulation isomorphism via IsoSigs, we select a variety of 3-manifolds to focus on, in particular those which span a range of interesting properties. Eight 3-manifolds were selected for these investigations, they include: the 3-sphere $S^{3}$; the product of the 2-sphere and the 1-sphere $S^{2}\times S^{1}$; 3-dimensional real projective space $\mathbb{RP}^{3}$, the lens space $L(7,1)$, the lens space $L(7,2)$; the 3-torus $T^{3}$; the Poincaré homology sphere $PHS$; and the smallest closed hyperbolic manifold $H_{SC}$. The 3-manifold abbreviations, along with the initial IsoSigs used to generate the datasets of equivalent triangulation IsoSigs are given in Table 1. The IsoSig encoding is of particular interest due to its exceptionally compressed nature, and a thorough introduction to its structure is given in Appendix A. Of these manifolds $S^{2}\times S^{1}$ is the only one not admitting a constant curvature isotropic complete Riemannian metric (we shall say for short a ‘non-isotropic manifold’), whilst the others are split amongst their curvatures, with $(S^{3},\mathbb{RP}^{3},L(7,1),L(7,2),PHS)$ positive curvature, $T^{3}$ zero curvature, and $H_{SC}$ negative curvature. In addition they have properties as listed in Table 2.
This selection of manifolds is somewhat arbitrary, but includes a list of the prototypical manifolds often considered in other works. Specifically $S^{3}$, $T^{3}$, represent the heuristically ‘simplest’ manifolds with $\{+,0\}$ curvatures, while $H_{SC}$ is selected as a representative with -- curvature. In fact, this choice is the hyperbolic 3-manifold with minimum volume (given in Table 2, contrasting to the other selected manifolds which all have 0 volume). $S^{2}\times S^{1}$ is a common choice for considering composite manifolds which may be considered as a direct product of lower dimensional manifolds. $\mathbb{RP}^{3}$ is the simplest 3d projective space, a particular important construction with a broad range of uses. Then, $PHS$ is the only homology 3-sphere with finite fundamental group (other than $S^{3}$ itself), making this an interesting comparison case for differentiating with $S^{3}$. Finally the two Lens spaces were the first constructions of 3-manifolds which were not differentiable by their homology and fundamental group alone (Table 2 shows that these properties match). Under special (more trivial) conditions of the Lens space constructions both $S^{3}$ and $S^{2}\times S^{1}$ can be produced, making these four a particularly interesting and potentially difficult set to differentiate between.
@@
For each 3-manifold’s initial triangulation and IsoSig, a dataset was generated of equivalent triangulations (each represented with their respective IsoSig) from the manifold’s Pachner graph computed exhaustively up to as high a depth (i.e. integer number of moves away from the initial IsoSig) as feasible within memory limits. The Pachner graphs were generated with moves $\{2-3,3-2\}$ (with further functionality in the code to consider $\{1-4,4-1\}$ moves). Preliminary investigations by these authors for this work using supervised neural networks to classify 3-manifolds from IsoSig triangulations showed promising performances with accuracies $>0.99$ for near-perfect learning. However, these architectures relied heavily on the IsoSig length for distinguishing manifolds; where gradient saliency methods revealed that the only used input was the first padded entry indicating that change of IsoSig input length was the dominantly used feature in the classification.
Since the IsoSig length is heavily correlated with the number of tetrahedra, when we initiate generation using triangulations with a similar (e.g. small) number of tetrahedra then the graphs will display the same IsoSig lengths throughout, however the frequency distribution of the observed lengths varies, and this is what the architectures can leverage to provide efficient estimates of the considered manifold (note the final depth often all have the same length and dominate these distributions). For the manifolds considered the IsoSig lengths observed in the deep Pachner graph generation were: $(9,11,14,17,19,22,25,27,30,33,35,38,41)$. The respective distributions are shown in Figure 2.
@@ 3.1.1 Orientable Cusped Census
Figure 6: The distribution of the sizes of the Pachner graphs (i.e. number of nodes) as a function of (a) the volume and of (b) the length of the shortest geodesic (systole). In blue the points corresponding to one cusped manifolds, in red those corresponding to $2$-cusped manifolds, and in green the remaining ones. Plot (c) shows plot (b) with axes logged to inform on the power law behaviour of the lower bound, including a potential lower bound of $\frac{75}{\sqrt{\ell}}$.
@@
We also analysed the size of the Pachner graphs as a function of the hyperbolic volume of the manifolds and their number of cusps (which varies from $1$ to $3$ in our dataset). The result is shown in Figure 6(a). Looking at the figure one realises that there are ‘peaks’ in the sizes of the Pachner graphs for one cusped manifolds and that these peaks correspond to the first red dots, namely the lowest volume $2$-cusped manifolds. This is no random phenomenon. The blue dot with highest size of the Pachner graph and volume $<3.63$ turns out to be “v0117” (in snappy’s nomenclature, IsoSig: hLALPkbcbefggghxwnxibk) and drilling out its shortest geodesic one gets “m129” (IsoSig: eLPkbdcddhgggb) which is the second lowest volume manifold in snappy’s census with number of cusps $2$ (it corresponds to one of the two leftmost red dots in the figure). This leads us to think that the size of the Pachner graph could be inversely related to the length of the shortest geodesic in the hyperbolic manifold which, by Thurston-Jorgensen’s theorem on hyperbolic Dehn fillings corresponds to Dehn fillings of manifolds with one more cusp with very large filling slopes. In order to check this we plotted in Figure 6(b), the sizes of the Pachner graphs with respect to the length $\ell$ of the shortest geodesic in the cusped manifold and the plot clearly shows that such number is bounded below by an unknown decreasing function of $\ell$. Scaling the plot axes, as in Figure 6(c), shows the behaviour more clearly; the lower bound forms a straight line indicating a power law relationship, and the gradient is $\sim-0.5$, indicating a relationship of the form $\propto\frac{1}{\sqrt{\ell}}$ for this dataset.
Furthermore, a closer look reveals that the dots come in families, corresponding to which manifold is obtained by drilling the shortest geodesic. This is exemplified in Figures 7(a) and 7(b) where the same analysis as before is restricted to the manifolds whose drilling is a fixed one, here “m129” (chosen just as an example). All this seems to corroborate the following
@@
The growth rate of the Pachner graph of a cusped hyperbolic $3$-manifold $M$ is bounded below by a function of the inverse of the systole (i.e. the length of the shortest geodesic in $M$).
One may be concerned by a potential bias in the previous analysis: as the number of tetrahedra grows the relative scarcity of triangulations with big shortest geodesic length also increases, and this may be causing the observed phenomenon. To check this, we first drew the histogram of the shortest geodesic lengths for the manifolds with complexity up to 8 in the Orientable Cusped Census of snappy: see Figure 8(a). The figure clearly shows that triangulations with shortest geodesic $<0.2$ are fairly common, thus in the right part of the same figure we restrict the image of Figure 7(b) to the interval $\ell\in[0,0.2]$. The results are consistent with the observed behaviour.
(a) Histogram of Shortest Geodesic Length
(b) Zoom in on Systoles $<0.2$
Figure 8: On the left the global histogram of the shortest geodesic length on the Oriented Cusped Census of snappy with $<9$ tetrahedra. On the right the restriction of the plot of the size of the Pachner graph for manifolds with shortest geodesic less than $0.2$. In green, the graph of the function $f(\ell)=\frac{75}{\sqrt{\ell}}$ seems to approximately minor the scatter plot on the interval $[0,0.2]$, but it is plot here only as an example.
@@ 4.1 Differentiating Manifolds
The NN architecture is the prototypical machine learning model, it is very adaptable, and used ubiquitously with great success throughout many fields. Any NN is made up of sets of neurons organised into layers, where each neuron takes in a vector of inputs, and outputs a single number. The neuron acts on the input vector x with a linear function, then a non-linear ‘activation’ function, such that: $\textbf{x}\mapsto act(\textbf{w}\cdot\textbf{x}+b)$ for the trainable parameters of weights vector w and bias scalar $b$. The full NN function of $L$-layers then takes the form
$f_{NN}:=\text{softmax}(W_{i_{L},i_{L-1}}\cdot\alpha(\ ...\ \alpha(W_{i_{1},i_{input}}\cdot x_{input}+b_{i_{1}})\ ...\ )+b_{i_{L}})\;.$ (4.1)
Demonstrating the vectorised linear action of matrix multiplication with the weight matrix $W_{i_{m}i_{n}}$ between layers $i_{m}$ and $i_{n}$, then addition of a bias vector $b_{i_{m}}$. Followed by the $\alpha$ activation function used to provide the alternating non-linear action; here we use the Leaky-ReLU activation function, which has significant speed advantages due to its simplicity, and effectively makes the training process the fitting of a piecewise linear function. Finally the softmax activation function is used to bound outputs into the range $[0,1]$ such that they produce a probability distribution for the classification task. These activation functions are defined
[…]
$\text{softmax}(x)_{i}$ $=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}\,,$ (4.3)
for some small factor $\varepsilon$ ($\sim 0.01$). These architectures are very adaptable, possess some universal approximation theorems supporting their use, and have practically been particularly successful across other fields of mathematics. More information on this architecture can be found at [76].
@@
Generally, transformer models are another class of supervised architectures designed to handle sequential language-like data using attention mechanisms instead of traditional recurrence or convolution. At their core, they employ an encoder-decoder structure, though many applications focus on encoder-only (e.g., BERT) or decoder-only (e.g., GPT) variants. The multi-head self-attention mechanism is the cornerstone, which allows the model to dynamically weigh the importance of all input tokens relative to each other. For an input sequence abcd... attention scores are computed using query $Q$, key $K$, ad value $V$ matrices, from which the attention importance weighting between letters of the input is computed as
$\text{Attention}(Q,K,V):=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V\;,$ (4.4)
for $d_{k}$ the dimensionality of the key vectors. This mechanism enables Transformers to capture long-range dependencies efficiently.
Transformers work via alternating small NNs providing trainable parameters for the encoding and decoding, and attention layers which preserve the interrelations within the sequence. Transformers are trained using a supervised loss function in a similar manner to NNs, typically using cross-entropy loss, over a labelled dataset. They are also particularly popular as the attention mechanism allows parallelisation in computation, making larger models more feasible than traditional recurrent methods. More information on this architecture can be found at [78].
As motivated in §2.2, the task of identifying equivalent triangulations is infeasibly expensive, and rigorous algorithm design is impractical. Therefore efficient statistical methods may prove useful in informing on the likelihood of equivalence, thereby practically informing on optimal resource allocation for more rigorous checks; i.e. if a statistical method says a pair IsoSigs are very likely to be equivalent, it may be worth committing resources to compute the connecting path where it would be les likely found for other pairs. Recent years have seen ML dominate the development of computational statistical methods across fields, and in this section we provide the first investigation into its efficacy on this equivalence problem and for general 3-manifolds as a whole. In doing so we consider [^10] More specialised architectures may well perform better, but we leave this experimentation to future work, since here we are performing the first tests of these methods for this problem. the ubiquitous NN architecture, and the transformer architecture which has shown particular success on language-like data (alike our IsoSig encodings).
@@ 4.1.1 Neural Networks
We performed an in-depth analysis of the gradient saliency values used by our NNs to distinguish $L(7,1)$ vs $L(7,2)$. A trained NN represents a function from inputs to outputs, in the case of these problems this is an input of a single IsoSig, one-hot encoded from length 30 in a length 64 alphabet to a length 1920 vector and an output of a length 2 vector with each entry evaluating in the range $[0,1]$ (as a probability distribution over the 2 classes) whose larger value indicates the predicted manifold the input IsoSig triangulation represents.
@@
Further examining the actual saliency values (represented by colours in Figure 13), averaged over all of the cross-validation runs as well as each runs test dataset, we can plot a histogram of the average saliency size, as shown in Figure 14(a). This image shows a partition of values with a selection having normalised absolute gradients $>10^{-4}$ and the majority having values $<10^{-6}$. There is therefore a clear selection of ‘important’ inputs and ‘non-important’ ones. Taking the input indices corresponding to the larger ‘important’ saliency values $>10^{-4}$, these can be sorted into what character in the alphabet they correspond to (as plotted in a histogram in Figure 14(b)), or into what character in the length 30 IsoSig they correspond to (as plotted in a histogram in Figure 14(c)).
@@
Figure 14(b) shows that the 363 ‘important’ inputs correspond to only a subset of the full alphabet – predominantly the beginning of this alphabet with a few letters deeper in. Equivalently, these same important input features have much higher frequencies towards the end of the IsoSig as shown in Figure 14(c). From this saliency analysis the NNs are hence relying most on the final entries of the IsoSigs (the permutation sequence) to determine the manifold, and the value of these entries becomes discriminant. Additionally the occurrence of specific letters (like h or i with the highest frequency in Figure 14(b)) can significantly help the architectures’ confidence in classification.